OpenStudy (anonymous):

5x-2y=20 solve for y. I got y=-2+x

OpenStudy (anonymous):

I got y=5/2x-10

OpenStudy (anonymous):

i also got what @kaylific did

OpenStudy (anonymous):

Can you go step by step though that? I have no idea what I did wrong?

OpenStudy (anonymous):

2y=5x-20 y=x(5/2)-20/2 y=5/2(x)-10

OpenStudy (anonymous):



OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

U wanna start by getting y on one side by itself. Do that by adding 2y and subtracting 20